OpenStudy (anonymous):

let f(x) =7x^2 -8 Find f(a^2)

OpenStudy (anonymous):

substitute a^2 in x's place. f(a^2)=7(a^2)^2-8 Simplify this.

OpenStudy (anonymous):

$$ f(a^2) = 7(a^2)^2- 8 = 7a^4 - 8 $$